40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 16:29:40 2020, gautam, Update, Electronics, c1psl VME crate removed, Acro-crate installed c1psl.pdf
    Reply  Tue Mar 3 17:20:14 2020, gautam, Update, Electronics, More cabling removed 
       Reply  Wed Mar 4 16:18:31 2020, gautam, Update, Electronics, More cabling removed 1X2EuroBefore.JPGIOO.png
          Reply  Wed Mar 4 21:02:49 2020, Koji, Update, Electronics, More cabling removed 
    Reply  Tue Mar 3 17:59:33 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
       Reply  Thu Mar 5 15:03:48 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
          Reply  Thu Mar 5 19:45:23 2020, Jon, Summary, PSL, C1PSL in-situ test results c1psl_feedthrough_wiring_-_By_Connector_(3).pdf
             Reply  Wed Mar 11 18:12:53 2020, gautam, Summary, PSL, WFS Demod board modifications D980233-B_Mar2020Mods.pdf
Message ID: 15266     Entry time: Wed Mar 11 18:12:53 2020     In reply to: 15256
Author: gautam 
Type: Summary 
Category: PSL 
Subject: WFS Demod board modifications 

[koji, gautam]

Attachment #1 shows the relevant parts of the schematic of the WFS demod board (not whitening board). 

  • The basic problem was that the switchable gain channels were not accounted for in the Acromag channel list 😒.
  • What this meant was that the DC gain was set to the default x100 (since the two DG211s that provide the switchable x10 and x1 gain options had their control logic pins pulled up to +5V because these pins weren't connected to any sinking BIO channel).
  • Rather than set up new connections to Acromags inside the chassis (though we have plenty of spares), Koji and I decided to make these fixed to x1 gain.
  • The actual fix was implemented as shown in the annotated schematic. There are some pictures 📷 of the PCB in the DCC entry linked above.
  • Amusingly, this board will now require a sourcing BIO unit if we want to still have the capability of switching gains.

Before removing the boards from the eurocrate: 

  • I dialled down the Kepco HV supplies
  • disconnected all the cabling to these boards after noting down cable numbers etc.

After Koji effected the fix, the boards were re-installed, HV supplies were dialled back up to nominal voltage/currents, and the PMC/IMC were re-locked. The WFS DC channels now no longer saturate even when the IMC is unlocked 👏 👏 . I leave it to Yehonathan / Jon to calibrate these EPICS channels into physical units of mW of power. We should also fix the MEDM screen and remove the un-necessary EPICS channels.

Later in the evening, I took advantage of the non-saturated readbacks to center the beams better on the WFS heads. Then, with the WFS servos disabled, I manually aligned the IMC mirrors till REFLDC was minimized. Then I centered the beam on the MC2 transmission QPD (looking at individual quadrants), and set the WFS1/2 RF offsets and MC2 Trans QPD offsets in this condition.


WFS DC channels are saturating when the IMC is unlocked.

Attachment 1: D980233-B_Mar2020Mods.pdf  223 kB  Uploaded Thu Mar 12 01:35:52 2020  | Show | Show all
ELOG V3.1.3-